# Nash equilibrium seeking for a class of quadratic-bilinear Wasserstein distributionally robust games

### v1

@@ 2.2 Reformulation as a data-driven variational inequality problem
An advantage of this reformulation is that it has good scalability properties with respect to the data size, which is important for data-driven applications. Equilibrium seeking using available algorithms in the literature often leads to a large number of oscillations, which are avoided with our problem formulation. Thus, through Theorem 1, we can obtain data-scalable reformulations for the class of heterogeneous data-driven Wasserstein distributionally robust games in (1). In the next section, we assess the computational performance of our theoretical results through an illustrative example and a risk-aware portfolio allocation game, which takes into account behavioural coupling of the investors’ decisions.
@@ 3 Numerical simulations
In the simulation results, we use two algorithms to solve the variational inequality problems: the adaptive golden ratio algorithm (aGRAAL) malitsky_golden_2020 , and the Hybrid method, Algorithm 1 (Hybrid-Alg). The Hybrid method is similar to aGRAAL but differs in the choice of the momentum parameter, as will be explained below.
3.1 Illustrative example
In this section, we reformulate a case study of the distributionally robust game in (1), under Assumption 1, as a variational inequality problem and solve it using both aGRAAL and Hybrid-Alg. The key difference between these two algorithms is that, unlike aGRAAL, which uses a fixed momentum parameter, Hybrid-Alg employs a variable momentum parameter. We believe this is a testimony to the potential of switching the momentum parameter between a small (used in aGRAAL) and a large value, which has a significant impact on convergence speed. In particular, having larger, variable momentum parameter in Algorithm 1 makes $\bar{x}_{k}$ closer to the most recent iterate, $x_{k}$, rather than $\bar{x}_{k-1}$, which allows us to estimate the local Lipschitz constant of the corresponding VI mapping $F$ more precisely compared to aGRAAL.
@@ 3.1 Illustrative example
6:  compute the following summations with $\phi_{k+1}=\bar{\phi}$:      $\text{sum}_{k+1}^{1}=\text{sum}_{k}^{1}$ + (Eq. 16 in Reza_2024 )       $\text{sum}_{k+1}^{2}=\text{sum}_{k}^{2}$ + (Eq. 17 in Reza_2024 )
@@
16:        $\text{sum}_{k+1}^{2}=\text{sum}_{k}^{2}$ + (Eq. 17 in Reza_2024 with $\phi_{k+1}=\alpha$)

### v2

@@ 3.2 Challenges of distributionally robust heterogeneous games compared to DRO
Even though the mapping can in general be nonmonotone, we illustrate how equilibria can still be computed efficiently based both on the structure of our problem, as provided by Theorem 1, and the equilibrium seeking algorithms we propose in the next section. Specifically, leveraging the structure of our reformulation, we avoid introducing epigraphic variables that render both the decision and the constraints dependent on the size of the data netessine_wasserstein_2019 , mohajerin_esfahani_data-driven_2018 . Thus, through Theorem 1, we can obtain reformulations for the class of heterogeneous data-driven Wasserstein distributionally robust games in (1) that scale better with data compared to the use of epigraph forms. In the next section, we assess the computational performance of our theoretical results through an illustrative example and a risk-aware portfolio allocation game, which takes into account behavioural coupling of the investors’ decisions.
@@ 4 Numerical Simulations
6:  compute the following summations with $\phi_{k+1}=\bar{\phi}$:      $\text{sum}_{k+1}^{1}=\text{sum}_{k}^{1}$ + (4.1)      $\text{sum}_{k+1}^{2}=\text{sum}_{k}^{2}$ + (22)
@@
16:        $\text{sum}_{k+1}^{2}=\text{sum}_{k}^{2}$ + ((22) with $\phi_{k+1}=\alpha$)
@@
In the simulation results, we use three algorithms to solve the variational inequality problems:
  1. (i)
    Forward-backward splitting (FB):
    $x^{k+1}=\text{prox}_{\tau g}(x^{k}-\tau F(x^{k})),$
    where $\tau$ is the stepsize. This method is known as a projected gradient descent algorithm as well which is the most used method in machine learning and control applications. The convergence of this method is guaranteed for strongly monotone (with a strongly monotone constant $\mu$) and Lipschitz (with a Lipschitz constant $L$) operator with $\tau\in\left(0,2\mu/L^{2}\right)$ (Pang1, , Theorem 12.4.6).
  2. (ii)
    Adaptive golden ration algorithm (aGRAAL) malitsky_golden_2020 :
    $\phi$ $\in(1,\frac{1+\sqrt{5}}{2}],\quad\rho=\dfrac{1}{\phi}+\dfrac{1}{\phi^{2}}$
    $\tau_{k}$ $=\min\left\{\rho\tau_{k-1},\dfrac{\phi\theta_{k-1}}{4\tau_{k-1}}\dfrac{\|x^{k}-x^{k-1}\|^{2}}{\|F(x^{k})-F(x^{k-1})\|^{2}},\bar{\tau}\right\}$
    $\bar{x}^{k}$ $=\dfrac{(\phi-1)x^{k}+\bar{x}^{k-1}}{\phi}$
    $x^{k+1}$ $=\text{prox}_{\tau_{k}g}(\bar{x}^{k}-\tau_{k}F(x^{k})),\quad\theta_{k+1}=\dfrac{\phi\tau_{k}}{\tau_{k-1}}$
    The convergence of this method is guaranteed for Lipschitz and monotone operator (with a Lipschitz constant $L$) (malitsky_golden_2020, , Lemma 3).
  3. (iii)
    Hybrid method (Hybrid-Alg): The full steps of this method can be found in Algorithm 1. This algorithm is presented in the recent work Reza_2024 and includes some of the authors of the present paper. However, we note that the convergence rate proof only holds for monotone mappings and is applied on a significantly less complicated case study. The application of this algorithm on our problem shows that the equilibria of even a generally nonmonotone mapping could be numerically obtained relatively fast, even faster than the golden ratio method in practice. The Hybrid method, though similar to aGRAAL, differs significantly in the choice of the momentum parameter, as we explain later on.
@@
4.1 Numerical Example
In this section, we reformulate a case study of the distributionally robust game in (1), under Assumption 1, as a variational inequality problem and solve it using FB (with $\tau=0.001$), aGRAAL, and Hybrid-Alg. The key difference between aGRAAL and Hybrid-Alg algorithms is that, unlike aGRAAL, which uses a fixed momentum parameter, Hybrid-Alg employs a variable momentum parameter. We believe this is a testimony to the potential of switching the momentum parameter between a small (used in aGRAAL) and a large value, which has a significant impact on convergence speed. In particular, having larger, variable momentum parameter in Algorithm 1 makes $\bar{x}_{k}$ closer to the most recent iterate, $x_{k}$, rather than $\bar{x}_{k-1}$, which allows us to estimate the local Lipschitz constant of the corresponding VI mapping $F$ more precisely compared to aGRAAL.
Briefly speaking, the following two equations are evaluated in each iteration of Hybrid-Alg (Algorithm 1) to assess sufficient decrease and to determine whether the large or the small momentum parameter should be used.
$\frac{\theta_{k-1}}{2}\|x^{k}-x^{k-1}\|^{2}$ $-\dfrac{\tau_{k}}{\tau_{k-1}}\phi_{k}\|x^{k}-\bar{x}^{k}\|^{2}+\bigl{(}\dfrac{\tau_{k}}{\tau_{k-1}}\phi_{k}-1-\frac{1}{\phi_{k+1}}\bigr{)}\|x^{k+1}-\bar{x}^{k}\|^{2}$
$-\bigl{(}\dfrac{\tau_{k}}{\tau_{k-1}}\phi_{k}-\theta_{k}\bigr{)}\|x^{k+1}-x^{k}\|^{2}-\frac{\theta_{k}}{2}\|x^{k+1}-x^{k}\|^{2}.$ (21)
$\color[rgb]{0,0,0}-\dfrac{\tau_{k}\phi_{k}}{\tau_{k-1}}\|x^{k}-\bar{x}^{k}\|^{2}+\bigl{(}\dfrac{\tau_{k}\phi_{k}}{\tau_{k-1}}-1-\frac{1}{\phi_{k+1}}\bigr{)}\|x^{k+1}-\bar{x}^{k}\|^{2}-\bigl{(}\dfrac{\tau_{k}\phi_{k}}{\tau_{k-1}}-\theta_{k}\bigr{)}\|x^{k+1}-x^{k}\|^{2}.$ (22)